Two similar companies had been competing since they first opened for business in the same year, 2013. At the end of 2016, Elemental Corporation reported assets of $42,500 and Stockholders Equity of $17,325. Blackman Enterprises reported assets of $125,000 and liabilities of $53,500.

a. Which company had the greater reported liabilities, Elemental or Blackman?
b. Which company provides a better investment opportunity? Why?
Answer #1

Solution:-

(a)

Reported liabilities (Elemental)= Assets-Stockholders equity = $42,500-$17,325 = $25,175

Reported liabilities (Blackman)= $53,500

As we can see from above, Blackman has greater reported liabilities of $53,500 as compared to $25,175 of Elemental.

(b)

Debt to total assets ratio (Elemental)= Outside Liabilities/Assets = 25,175/42,500 = 59.24%

Debt to total assets ratio (Blackman)= Outside Liabilities/Assets = 53,500/125,000 = 42.8%

Blackman has a lower debt to total assets ratio which means it carries a smaller financing risk. Moreover, it has higher total assets which means it has a business of a larger size as compared to Elemental which means higher market share in the industry. Due to these factors, Blackman offers a better investment opportunity.